John is 4 times as old as Martha. Eight years ago the sum of their ages was sixty-nine. How old are they now?

Answer by jorel1380(3719)

J=4M
(J-8)+(M-8)=69
-----------
J+M-16=69
4M+M=85
5M=85
M=17
J=68
John is 68. Martha is 17..